OpenStudy (anonymous):

f(x)=1-x/3x+2. Find f^-1(x)

OpenStudy (anonymous):

\[x=\frac{1-y}{3y+2}\] solve for \(y\) it takes a few steps

OpenStudy (anonymous):

Can you type out the steps?

OpenStudy (anonymous):

1) multiply both sides by \(3y+2\) to get \[x(3y+2)=1-y\]

OpenStudy (anonymous):

2) distribute on the left to get \[3xy+2x=1-y\]

OpenStudy (anonymous):

3) put all terms with \(y\) on one side, all else on the other \[3xy+y=1-2x\]

OpenStudy (anonymous):

4) factor out the \(y\) on the left \[(3x+1)y=1-2x\]

OpenStudy (anonymous):

5) divide both sides by \(3x+1\) \[y=\frac{1-2x}{3x+1}\]

OpenStudy (anonymous):

6) write \[f^{-1}(x)=\frac{1-2x}{3x+1}\]

OpenStudy (anonymous):

THANKS!

OpenStudy (anonymous):

it was a lot of steps check my algebra, but i think it is right yw

OpenStudy (anonymous):

It's correct.